(Bond valuation) Enterprise, Inc. bonds have an annual coupon rate of 11 percent. The interest is paid semiannually and the bonds mature in 9 years. Their par value is $1,000. If the market's required yield to maturity on a comparable-risk bond is 14 percent, what is the value of the bond? What is its value if the interest is paid annually and semiannually? (Round to the nearest cent.)


 a. The value of the Enterprise bonds if the interest is paid semiannually is _______ 

 b. The value of the Enterprise bonds if the interest is paid annually is _______ 
Answer #1

a:

In case of semi annual payments

Using financial calculator
Input:

FV= 1000,

PMT=11%*1000/2 = 55,

N=9*2 = 18

I/Y =14/2 = 7

Find PV = $ 849.11

b: Annual payments

Using financial calculator
Input:

FV= 1000,

PMT=11%*1000 = 110,

N=9

I/Y =14

Find PV = $ 851.61